# Fully Dynamic Euclidean $k$-Means

### v2

@@ 11 Restricted $k$-Means
$\textsc{cost}(S,S-R^{*})\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*})+\sum_{s\in R^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}).$
@@
$\sum_{s\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{x\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}),$ (30)
$\sum_{s\in\widehat{R}^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\sum_{s\in R^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}).$ (31)
@@
$\mathrm{dist}(s,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(s,s^{\prime})\leavevmode\nobreak\ =\leavevmode\nobreak\ \mathrm{dist}(s,S-R^{*}).$
@@
$\forall i\in[t]:$ $w_{S}({\widehat{r}}^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}(r^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}(r^{*}_{i},S-r^{*}_{i}).$
@@
$w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}({\widehat{r}}^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}(r^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}(r^{*}_{i},S-r^{*}_{i})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-r^{*}_{i})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-R^{*}).$
@@
$w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-\widehat{R}^{*})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},s^{\prime})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 4\cdot w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-R^{*}).$
@@
Clearly, the above process will remove at most $2\cdot|R^{*}|=2r$ centers from $T_{1}$, thus leaving at least $6r-2r=4r$ centers still in $T_{1}$ (1 of Algorithm 4).
@@
[…]
Further, for every $s\in T^{\prime}_{1}$, it holds that $g(s)\in g(T_{1})-R^{*}=S-R^{*}$ and $\mathrm{dist}(s,g(s))=\mathrm{dist}(s,S-s)$. We then apply the following claim from [Tho04].
Claim 11.6 ([Tho04, Lemma 21]).
@@
$T^{\prime}_{1}\leavevmode\nobreak\ :=\leavevmode\nobreak\ \bigcup_{i\in[t]}(g^{-1}(g_{i})-s^{\prime}_{i}).$ (32)
@@
[…]
@@
$\textsc{cost}(T,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
@@
$\textsc{cost}(T,T-R)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-R)$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),\ \mathrm{dist}^{2}(s,T_{2})\}.$
@@
$\textsc{cost}(T,T-R)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),\mathrm{dist}^{2}(s,T_{2})\}$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),O(\epsilon^{-3})\cdot\mathrm{dist}^{2}(s,S-T_{1})\}$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\sum_{s\in R}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R)$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
@@ 12 Augmented $k$-Means
[…]
2:Sample $t:=\Theta(\epsilon^{-6}\cdot d\log\Delta)$ i.i.d. points $\{x_{i}\}_{i\in[t]}$ from $X$, such that $\Pr[x_{i}=y]\geq\Omega(\epsilon^{6})\cdot\frac{w(y)\cdot\mathrm{dist}^{2}(y,S+A)}{\textsc{cost}(X,S+A)}$ for every $y\in X$.
[…]
@@
We modify Algorithm 5 as follows: in each of the $a+1$ iterations, instead of only adding the points $\{x_{i}\}_{i\in[t]}$ (sampled in 2) to the set of augmented centers $A$, we also add them to the center set $S$ and update the data structure of Corollary 10.16. This update takes a total of $\tilde{O}(2^{\epsilon d})$ time per iteration. With this modification, 2 becomes equivalent to sampling points $x\in X$ with probability at least $\Omega(\epsilon^{6})\cdot\frac{w(x)\cdot\mathrm{dist}(x,S)}{\textsc{cost}(X,S)}$. By querying the data structure of Corollary 10.16 $t$ times, we obtain $t$ independent samples in total time $\tilde{O}(2^{\epsilon d})\cdot t=\tilde{O}(2^{\epsilon d})$. Hence, the total running time for the entire for loop is $\tilde{O}(2^{\epsilon d}\cdot a)$. After the for loop completes, we perform an additional step: we remove the augmented centers $A$ from $S$ and update the data structure from Corollary 10.16 accordingly. This step takes $|A|\cdot\tilde{O}(2^{\epsilon d})=\tilde{O}(2^{\epsilon d}\cdot a)$ time.
@@
The proof of Claim 12.3 follows almost directly from [ADK09, Theorem 1], where a similar sampling procedure is analyzed to obtain a bi-criteria approximation to the vanilla clustering problem. For completeness, we provide a proof in Section 12.1.
@@
A crucial step in the analysis is to understand the behavior of the (distorted) $D^{2}$-sampling procedure, as described in 2 of Algorithm 5 – namely, sampling a random point $x\in X$ such that
@@
$\textsc{cost}(X,S+A)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \textsc{cost}(X_{0},S+A)+\sum_{j=1}^{a}\textsc{cost}(X_{j},S+A)$
$\leq\leavevmode\nobreak\ \textsc{cost}(X_{0},S+A^{*})+\sum_{j=1}^{a}8\cdot\textsc{cost}(X_{j},S+A^{*})$
$\leq\leavevmode\nobreak\ 8\cdot\textsc{cost}(X,S+A^{*}).$
@@
$\widehat{X}_{j}\leavevmode\nobreak\ :=\leavevmode\nobreak\ \left\{y\in X_{j}:\ \mathrm{dist}^{2}(y,c_{j})\leq\frac{2}{w(X_{j})}\cdot\textsc{cost}(X_{j},c_{j})\right\}.$ (37)
A random point $x\in X$ sampled via $D^{2}$-sampling (2 of Algorithm 5) turns out to satisfy:
@@
    Pr[xX^j for some $\Pr[\text{$x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$}]\geq\Omega(\epsilon^{6})$.
@@
Item (a). According to the $D^{2}$-sampling procedure (2 of Algorithm 5), we have
Pr[xX^j for some $\Pr[\text{$x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$}]$ $\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \Omega(\epsilon^{6})\cdot\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\textsc{cost}(X,S+A)}$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \Omega(\epsilon^{6})\cdot\underbrace{\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}{\textsc{cost}(X,S+A)}\Bigg{.}}_{(\dagger)}\cdot\underbrace{\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}\Bigg{.}}_{(\ddagger)}.$
@@
For the term ()($\dagger$)( † ), we deduce that
(($\dagger$)( † ) = $\leavevmode\nobreak\ =\leavevmode\nobreak\ 1-\frac{\textsc{cost}(X_{0},S+A)+\sum_{j\neq 0\land j\notin\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}{\textsc{cost}(X,S+A)}$
$\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 1-\frac{\textsc{cost}(X_{0},S+A^{*})+\sum_{j\neq 0\land j\notin\mathcal{B}(A)}8\cdot\textsc{cost}(X_{j},S+A^{*})}{16\cdot\textsc{cost}(X,S+A^{*})}$
$\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{1}{2}.$
@@
For the term ()($\ddagger$)( ‡ ), let us consider a specific bad cluster $j\in\mathcal{B}(A)$. Let $\alpha:=\sqrt{w(X_{j})}\cdot\mathrm{dist}(c_{j},S+A)$ and $\beta:=\sqrt{\textsc{cost}(X_{j},c_{j})}$ for brevity. We can upper-bound $\textsc{cost}(X_{j},S+A)$ using these parameters.
$\textsc{cost}(X_{j},S+A)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{y\in X_{j}}w(y)\cdot\big{(}\mathrm{dist}(y,c_{j})+\mathrm{dist}(c_{j},S+A)\big{)}^{2}$ (triangle inequality)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{y\in X_{j}}w(y)\cdot\big{(}2\cdot\mathrm{dist}^{2}(y,c_{j})+2\cdot\mathrm{dist}^{2}(c_{j},S+A)\big{)}$ (AM-GM inequality)
= $\leavevmode\nobreak\ =\leavevmode\nobreak\ 2\beta^{2}+2\alpha^{2}.$
@@
$\textsc{cost}(\widehat{X}_{j},S+A)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{y\in\widehat{X}_{j}}w(y)\cdot\mathrm{dist}^{2}(y,S+A)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{1}{2}\cdot(\alpha-\sqrt{2}\beta)^{2}.$
@@
$(\ddagger)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{\left(\alpha-\sqrt{2}\beta\right)^{2}}{4\beta^{2}+4\alpha^{2}}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{(\sqrt{3}-\sqrt{2})^{2}}{4+4\cdot(\sqrt{3})^{2}}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{5-2\sqrt{6}}{16}.$
@@
$\textsc{cost}(X_{j},x)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{y\in X_{j}}w(y)\cdot\big{(}2\cdot\mathrm{dist}^{2}(y,c_{j})+2\cdot\mathrm{dist}^{2}(x,c_{j})\big{)}$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2\cdot\textsc{cost}(X_{j},c_{j})+4\cdot\textsc{cost}(X_{j},c_{j})$ (Equation 37)
= $\leavevmode\nobreak\ =\leavevmode\nobreak\ 6\cdot\textsc{cost}(X_{j},S+A^{*}).$ (Equation 35)
@@
For every $i\in[a+1]$, let $A_{i}$ denote the center set $A$ after the $i$-th iteration of the for loop (1, 2 and 3) in Algorithm 5. Also, let $A_{0}:=\emptyset$. We define the following two events for every $i\in[a+1]$:
@@
$\Pr\big{[}\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i}\big{]}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \Pr\big{[}\neg\mathcal{F}_{i}\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}\leavevmode\nobreak\ =\leavevmode\nobreak\ \Pr\big{[}|\mathcal{B}(A_{i})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}.$
Recall that $A_{i}=A_{i-1}+\{x_{1},\dots,x_{t}\}$, where $x_{1},\dots,x_{t}$ are the independent points sampled in the $i$-th iteration by 2. According to Claim 12.6, we have
$\forall j\in[t],\quad\Pr\big{[}|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}\leq 1-\Omega(\epsilon^{6}).$
@@
$\Pr\big{[}\forall j\in[t],|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}\leq\left(1-\Omega(\epsilon^{6})\right)^{t}\leq\frac{1}{\mathrm{poly}(\Delta^{d})}.$

### v3

@@ 12 Restricted $k$-Means
$\textsc{cost}(S,S-R^{*})\penalty 10000\ =\penalty 10000\ \sum_{s\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*})+\sum_{s\in R^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}).$
@@
$\sum_{s\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ \sum_{x\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}),$ (31)
$\sum_{s\in\widehat{R}^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(S,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\sum_{s\in R^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}).$ (32)
@@
$\mathrm{dist}(s,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(s,s^{\prime})\penalty 10000\ =\penalty 10000\ \mathrm{dist}(s,S-R^{*}).$
@@
$\forall i\in[t]:$ $w_{S}({\widehat{r}}^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})\penalty 10000\ \leq\penalty 10000\ w_{S}(r^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}(r^{*}_{i},S-r^{*}_{i}).$
@@
$w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$ $\penalty 10000\ \leq\penalty 10000\ w_{S}({\widehat{r}}^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$
$\penalty 10000\ \leq\penalty 10000\ w_{S}(r^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}(r^{*}_{i},S-r^{*}_{i})$
$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-r^{*}_{i})$
$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-R^{*}).$
@@
$w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-\widehat{R}^{*})$ $\penalty 10000\ \leq\penalty 10000\ w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},s^{\prime})$
$\penalty 10000\ \leq\penalty 10000\ 4\cdot w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$
$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-R^{*}).$
@@
Clearly, the above process will remove at most $2\cdot|R^{*}|=2r$ centers from $T_{1}$, thus leaving at least 6r2r=4r6r-2r=4r centers still in $T_{1}$ (1 of Algorithm 4).
@@
$|T^{\prime}_{1}|\penalty 10000\ =\penalty 10000\ \sum_{y\in g(T_{1})-R^{*}}|g^{-1}(y)|\penalty 10000\ \geq\penalty 10000\ |g(T_{1})-R^{*}|\penalty 10000\ \geq\penalty 10000\ |g(T_{1})|-|R^{*}|\penalty 10000\ \geq\penalty 10000\ 3r-r\penalty 10000\ =\penalty 10000\ 2r.$
Further, for every $s\in T^{\prime}_{1}$, it holds that $g(s)\in g(T_{1})-R^{*}=S-R^{*}$ and $\mathrm{dist}(s,g(s))=\mathrm{dist}(s,S-s)$. We then apply the following claim from [Thorup04].
Claim 12.6 ([Thorup04, Lemma 21]).
@@
$T^{\prime}_{1}\penalty 10000\ :=\penalty 10000\ \bigcup_{i\in[t]}(g^{-1}(g_{i})-s^{\prime}_{i}).$ (33)
@@
$\mathrm{dist}(s,s^{\prime}_{i})\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(s,g_{i})+\mathrm{dist}(s^{\prime}_{i},g_{i})\penalty 10000\ \leq\penalty 10000\ 2\cdot\mathrm{dist}(s,g_{i})\penalty 10000\ =\penalty 10000\ 2\cdot\mathrm{dist}(s,S-s).$
@@
$\textsc{cost}(T,T-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
@@
$\textsc{cost}(T,T-R)$ $\penalty 10000\ =\penalty 10000\ \sum_{s\in R}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-R)$
$\penalty 10000\ =\penalty 10000\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),\ \mathrm{dist}^{2}(s,T_{2})\}.$
@@
$\textsc{cost}(T,T-R)$ $\penalty 10000\ =\penalty 10000\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),\mathrm{dist}^{2}(s,T_{2})\}$
$\penalty 10000\ \leq\penalty 10000\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),O(\epsilon^{-3})\cdot\mathrm{dist}^{2}(s,S-T_{1})\}$
$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\sum_{s\in R}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R)$
$\penalty 10000\ =\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
@@ 13 Augmented $k$-Means
1: $A\leftarrow\emptyset$ $\triangleright$ The set of augmented centers
2: for all $i\in[a+1]$ do
3:   Sample $t:=\Theta(\epsilon^{-6}\cdot d\log\Delta)$ i.i.d. points $\{x_{i}\}_{i\in[t]}$ from $X$, such that $\Pr[x_{i}=y]\geq\Omega(\epsilon^{6})\cdot\frac{w(y)\cdot\mathrm{dist}^{2}(y,S+A)}{\textsc{cost}(X,S+A)}$ for every $y\in X$.
4:   $A\leftarrow A+\{x_{i}\}_{i\in[t]}$ return AA
@@
We modify Algorithm 5 as follows: in each of the $a+1$ iterations, instead of only adding the points $\{x_{i}\}_{i\in[t]}$ (sampled in 3) to the set of augmented centers $A$, we also add them to the center set $S$ and update the data structure of Corollary 11.16. This update takes a total of $\tilde{O}(2^{\epsilon d})$ time per iteration. With this modification, 3 becomes equivalent to sampling points $x\in X$ with probability at least $\Omega(\epsilon^{6})\cdot\frac{w(x)\cdot\mathrm{dist}(x,S)}{\textsc{cost}(X,S)}$. By querying the data structure of Corollary 11.16 $t$ times, we obtain $t$ independent samples in total time $\tilde{O}(2^{\epsilon d})\cdot t=\tilde{O}(2^{\epsilon d})$. Hence, the total running time for the entire for loop is $\tilde{O}(2^{\epsilon d}\cdot a)$. After the for loop completes, we perform an additional step: we remove the augmented centers $A$ from $S$ and update the data structure from Corollary 11.16 accordingly. This step takes $|A|\cdot\tilde{O}(2^{\epsilon d})=\tilde{O}(2^{\epsilon d}\cdot a)$ time.
@@
The proof of Claim 13.3 follows almost directly from [DBLP:conf/approx/AggarwalDK09, Theorem 1], where a similar sampling procedure is analyzed to obtain a bi-criteria approximation to the vanilla clustering problem. For completeness, we provide a proof in Section 13.1.
@@
A crucial step in the analysis is to understand the behavior of the (distorted) $D^{2}$-sampling procedure, as described in 3 of Algorithm 5 – namely, sampling a random point $x\in X$ such that
@@
$\textsc{cost}(X,S+A)\penalty 10000\$$=\penalty 10000\ \textsc{cost}(X_{0},S+A)+\sum_{j=1}^{a}\textsc{cost}(X_{j},S+A)$
$\leq\penalty 10000\ \textsc{cost}(X_{0},S+A^{*})+\sum_{j=1}^{a}8\cdot\textsc{cost}(X_{j},S+A^{*})$
$\leq\penalty 10000\ 8\cdot\textsc{cost}(X,S+A^{*}).$
@@
$\widehat{X}_{j}\penalty 10000\ :=\penalty 10000\ \left\{y\in X_{j}:\ \mathrm{dist}^{2}(y,c_{j})\leq\frac{2}{w(X_{j})}\cdot\textsc{cost}(X_{j},c_{j})\right\}.$ (38)
A random point $x\in X$ sampled via $D^{2}$-sampling (3 of Algorithm 5) turns out to satisfy:
@@
    $\Pr[\text{$x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$}]\geq\Omega(\epsilon^{6})$.
@@
Item (a). According to the $D^{2}$-sampling procedure (3 of Algorithm 5), we have
$\Pr[\text{$x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$}]$ $\penalty 10000\ \geq\penalty 10000\ \Omega(\epsilon^{6})\cdot\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\textsc{cost}(X,S+A)}$
$\penalty 10000\ =\penalty 10000\ \Omega(\epsilon^{6})\cdot\underbrace{\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}{\textsc{cost}(X,S+A)}\Bigg.}_{(\dagger)}\cdot\underbrace{\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}\Bigg.}_{(\ddagger)}.$
@@
For the term $(\dagger)$, we deduce that
$(\dagger)$ = $\penalty 10000\ =\penalty 10000\ 1-\frac{\textsc{cost}(X_{0},S+A)+\sum_{j\neq 0\land j\notin\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}{\textsc{cost}(X,S+A)}$
$\penalty 10000\ \geq\penalty 10000\ 1-\frac{\textsc{cost}(X_{0},S+A^{*})+\sum_{j\neq 0\land j\notin\mathcal{B}(A)}8\cdot\textsc{cost}(X_{j},S+A^{*})}{16\cdot\textsc{cost}(X,S+A^{*})}$
$\penalty 10000\ \geq\penalty 10000\ \frac{1}{2}.$
@@
For the term $(\ddagger)$, let us consider a specific bad cluster $j\in\mathcal{B}(A)$. Let $\alpha:=\sqrt{w(X_{j})}\cdot\mathrm{dist}(c_{j},S+A)$ and $\beta:=\sqrt{\textsc{cost}(X_{j},c_{j})}$ for brevity. We can upper-bound $\textsc{cost}(X_{j},S+A)$ using these parameters.
$\textsc{cost}(X_{j},S+A)$ $\penalty 10000\ \leq\penalty 10000\ \sum_{y\in X_{j}}w(y)\cdot\big(\mathrm{dist}(y,c_{j})+\mathrm{dist}(c_{j},S+A)\big)^{2}$ (triangle inequality)
$\penalty 10000\ \leq\penalty 10000\ \sum_{y\in X_{j}}w(y)\cdot\big(2\cdot\mathrm{dist}^{2}(y,c_{j})+2\cdot\mathrm{dist}^{2}(c_{j},S+A)\big)$ (AM-GM inequality)
= $\penalty 10000\ =\penalty 10000\ 2\beta^{2}+2\alpha^{2}.$
@@
$\textsc{cost}(\widehat{X}_{j},S+A)\penalty 10000\ =\penalty 10000\ \sum_{y\in\widehat{X}_{j}}w(y)\cdot\mathrm{dist}^{2}(y,S+A)\penalty 10000\ \geq\penalty 10000\ \frac{1}{2}\cdot(\alpha-\sqrt{2}\beta)^{2}.$
@@
$(\ddagger)\penalty 10000\ \geq\penalty 10000\ \frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}\penalty 10000\ \geq\penalty 10000\ \frac{\left(\alpha-\sqrt{2}\beta\right)^{2}}{4\beta^{2}+4\alpha^{2}}\penalty 10000\ \geq\penalty 10000\ \frac{(\sqrt{3}-\sqrt{2})^{2}}{4+4\cdot(\sqrt{3})^{2}}\penalty 10000\ =\penalty 10000\ \frac{5-2\sqrt{6}}{16}.$
@@
$\textsc{cost}(X_{j},x)$ $\penalty 10000\ \leq\penalty 10000\ \sum_{y\in X_{j}}w(y)\cdot\big(2\cdot\mathrm{dist}^{2}(y,c_{j})+2\cdot\mathrm{dist}^{2}(x,c_{j})\big)$
$\penalty 10000\ \leq\penalty 10000\ 2\cdot\textsc{cost}(X_{j},c_{j})+4\cdot\textsc{cost}(X_{j},c_{j})$ (Equation 38)
= $\penalty 10000\ =\penalty 10000\ 6\cdot\textsc{cost}(X_{j},S+A^{*}).$ (Equation 36)
@@
For every $i\in[a+1]$, let $A_{i}$ denote the center set $A$ after the $i$-th iteration of the for loop (2, 3 and 4) in Algorithm 5. Also, let $A_{0}:=\emptyset$. We define the following two events for every $i\in[a+1]$:
@@
$\Pr\big[\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i}\big]\penalty 10000\ \leq\penalty 10000\ \Pr\big[\neg\mathcal{F}_{i}\penalty 10000\ \big|\penalty 10000\ \neg\mathcal{E}_{i}\big]\penalty 10000\ =\penalty 10000\ \Pr\big[|\mathcal{B}(A_{i})|>|\mathcal{B}(A_{i-1})|-1\penalty 10000\ \big|\penalty 10000\ \neg\mathcal{E}_{i}\big].$
Recall that $A_{i}=A_{i-1}+\{x_{1},\dots,x_{t}\}$, where $x_{1},\dots,x_{t}$ are the independent points sampled in the $i$-th iteration by 3. According to Claim 13.6, we have
$\forall j\in[t],\quad\Pr\big[|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\penalty 10000\ \big|\penalty 10000\ \neg\mathcal{E}_{i}\big]\leq 1-\Omega(\epsilon^{6}).$
@@
$\Pr\big[\forall j\in[t],|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\penalty 10000\ \big|\penalty 10000\ \neg\mathcal{E}_{i}\big]\leq\left(1-\Omega(\epsilon^{6})\right)^{t}\leq\frac{1}{\mathrm{poly}(\Delta^{d})}.$